# Comparative Analysis of QSVM and VQC Models for B-cell Epitope Prediction in Vaccine Design Using Quantum Machine Learning

### v1

@@ Abstract
Support Vector Machine (SVM) is a classical and widely applied supervised machine learning algorithm for binary classification. Utilizing the kernel trick, it enables the embedding of data into higher-dimensional feature spaces to effectively address non-linear classification tasks.
This study primarily investigates the classical Support Vector Machine (SVM) and introduces an enhanced quantum version—Quantum Support Vector Machine (QSVM)—constructed using quantum circuits, to compare their performance on classification tasks. In QSVM, the high-dimensional computations typically performed by kernel methods are offloaded to quantum circuits, enabling data to be mapped into higher-dimensional feature spaces and potentially accelerating computation, with the aim of improving overall performance.
In addition, this study incorporates a comparison with the Variational Quantum Classifier (VQC), a purely quantum model that entirely replaces classical computation with quantum circuits. VQC is designed to harness the advantages of quantum computing in handling high-dimensional feature spaces, making it a promising approach for classification and prediction tasks.
In recent years, the outbreak of COVID-19 has underscored the importance of vaccine development. One critical step in this process is the prediction of B-cell epitopes. In this study, a B-cell epitope dataset from the Immune Epitope Database (IEDB) is used for model training and prediction. The dataset contains 10 features that describe each B-cell epitope.
To facilitate the analysis of model performance across different dimensions, Principal Component Analysis (PCA) is applied for dimensionality reduction. This allows us to examine how performance varies from lower- to higher-dimensional feature spaces. Additionally, the effect of sample size on prediction outcomes is investigated, treating it as a variable in the analysis.
QSVM and VQC were found to perform well under different conditions, with VQC showing stronger performance in larger datasets and QSVM being effective in low-sample regimes. This work provides new insights into the practical deployment of quantum machine learning for biomedical sequence classification tasks, especially in vaccine design scenarios.
@@ 1 Introduction
The original motivation behind quantum computing was to overcome the inherent limitations of classical computers. When simulating quantum systems on classical machines, the exponential growth of the Hilbert space leads to intractable computational costs. In 1982, Richard Feynman proposed using quantum systems to simulate quantum phenomena, laying the foundation for the field.
Quantum machine learning (QML), which integrates quantum computing with machine learning frameworks, has emerged as a promising field. Two notable quantum classifiers are the quantum support vector machine (QSVM) and the variational quantum classifier (VQC), both of which have shown potential in solving classification problems with high-dimensional data.
Recent studies have demonstrated the application of QSVM and VQC in biomedical domains. Zhuang et al. [25] successfully used a QSVM to classify non-hemolytic peptides, outperforming classical SVM models on biological datasets. Similarly, Li et al. [26] presented a comprehensive review of QML applications in drug discovery and immunoinformatics, reinforcing the role of quantum-enhanced models in medical research.
In this work, we investigate the use of QSVM and VQC for predicting linear B-cell epitopes — a crucial component of vaccine design. Using benchmark datasets from IEDB, we apply PCA-based preprocessing and conduct a detailed comparison of model performance under varying input dimensionality and sample sizes. Our results provide insight into the practical feasibility of QML approaches in computational immunology.
@@ 3 Background and Data Preprocessing
Figure 1: 2D PCA projection of the input features.
Figure 2: 3D PCA projection of the input features.
@@ 3.1 B-cell Epitope Prediction Background
B cells play a central role in the adaptive immune system by recognizing antigenic determinants (epitopes) on protein surfaces and producing antibodies specific to those regions. Accurate identification of linear B-cell epitopes is essential for rational vaccine design, as it enables the selection of high-immunogenicity protein segments for epitope-based vaccines.
To address the need for efficient and scalable epitope prediction, especially in the context of emergent viral threats such as SARS-CoV-2, immunoinformatics approaches have become increasingly prevalent. Databases such as the Immune Epitope Database (IEDB) and UniProt provide well-annotated repositories of peptide sequences and their associated immune activities. In our study, we focus on linear B-cell epitope data extracted from IEDB, specifically those associated with IgG-class antibodies, which represent the majority of curated entries.
@@ 4 Quantum Support Vector Machine, QSVM
Figure 3: Configuration diagram of the Quantum Support Vector Machine (QSVM).
@@ 4.1 Kernel-based Classification Using Quantum Circuits
$K(x_{i},x_{j})=|\langle\phi(x_{i})|\phi(x_{j})\rangle|^{2}$
The quantum feature maps $\phi(x)$ are implemented by parametrized quantum circuits. The resulting kernel matrix is estimated by repeatedly executing the circuit and measuring in the computational basis. The number of times the all-zero outcome is observed determines the kernel estimate. This quantum kernel is then passed to a classical SVM solver for final classification, forming a hybrid quantum-classical architecture.
@@ 4.2 Data Encoding and Feature Mapping
$|x\rangle=U_{\phi}(x)|0\rangle^{\otimes n}$
@@ 5 Variational Quantum Classifier, VQC
Figure 4: Configuration diagram of the Variational Quantum Classifier (VQC).
@@ 6.2 QSVM
Table 2 shows the results of the QSVM. The highest accuracy, 70%, was achieved when using 20 samples with 2-dimensional input features.
@@
Table 1: QSVM Accuracy
@@
Table 2: QSVM AUC value
@@
Table 3: QSVM MCC value
@@ 6.3 The Variational Quantum Classifier
Table 5 presents the accuracy of the VQC with 10 training epochs, where the sample size ranges from 1,000 to 4,000 and the input dimension varies from 2 to 10.
@@
Table 4: VQC 10 epoch Accuracy
As shown in Table 5, the accuracy generally hovers around 73%, with the highest accuracy achieved when the sample size is 1,000 and the input dimension is 4.
Table 6 presents the AUC values for 10 training epochs, with sample sizes ranging from 1,000 to 4,000 and input dimensions from 2 to 10.
@@
Table 5: VQC 10 epoch AUC value
As shown in Table 6, the best AUC value of 0.6750 was obtained when the sample size was 4,000 and the input dimension was 10.
Table 7 presents the MCC values for 10 training epochs, with sample sizes ranging from 1,000 to 4,000 and input dimensions from 2 to 10.
@@
Table 6: VQC 10 epoch MCC value
In Table 7, the highest MCC value of 0.0987 was achieved when the sample size was 3,000 and the input dimension was 4.
Table 8 presents the accuracy results for 100 training epochs, with sample sizes ranging from 1,000 to 4,000 and input dimensions from 2 to 10.
@@
Table 7: VQC 100 epoch Accuracy
The accuracy values in Table 8 are similar to those observed at 10 epochs, remaining around 73%. In this case, the highest accuracy was achieved with a sample size of 3,000 and an input dimension of 10.
Table 9 presents the AUC values for 100 training epochs, with sample sizes ranging from 1,000 to 4,000 and input dimensions from 2 to 10.
@@
Table 8: VQC 100 epoch AUC value
In Table 9, the AUC value reaches 0.7, which occurs at an input dimension of 10 with a sample size of 4,000—the largest sample size and the highest dimensional setting in the experiment.
Table 10 presents the MCC values for 100 training epochs, with sample sizes ranging from 1,000 to 4,000 and input dimensions from 2 to 10.
@@
Table 9: VQC 100 epoch MCC value
As shown in Table 10, the MCC value exceeds 0.1 when the sample size is 1,000 and the input dimension is 6, representing a relatively good performance.
Table 11 presents the accuracy results for 150 training epochs, with sample sizes ranging from 1,000 to 4,000 and input dimensions from 2 to 10.
@@
Table 10: VQC 150 epoch Accuracy
Compared to the previous two epoch settings (Tables 5 and 8), Table 11 shows a slight increase in accuracy, reaching nearly 74
Table 12 presents the AUC values for 150 training epochs, with sample sizes ranging from 1,000 to 4,000 and input dimensions from 2 to 10.
@@
Table 11: VQC 150 epoch AUC value
@@
Table 13 presents the MCC values for 150 training epochs, with sample sizes ranging from 1,000 to 4,000 and input dimensions from 2 to 10.
@@
Table 12: VQC 150 epoch MCC value
@@ 6.4 Effect of Feature Dimensionality
To investigate the influence of feature dimensionality on model performance, we conducted a series of experiments using Principal Component Analysis (PCA) to reduce input features to 2 through 10 components. Both QSVM and VQC were evaluated under these conditions using fixed training sizes (e.g., 1000, 2000, 4000 samples).
Our results indicate that model performance varies non-linearly with feature dimensionality. In particular, the VQC tends to perform better in low-to-medium dimensional regimes (e.g., 4–6 dimensions), whereas the QSVM is more stable across all tested dimensions. These observations suggest that VQC may benefit from feature compression under limited quantum hardware resources.
@@ 6.5 Comparison with Classical Models
To establish a baseline, we trained a classical Support Vector Machine (SVM) using the radial basis function (RBF) kernel on the same dataset and preprocessing pipeline. The classical SVM achieved an accuracy of approximately 70%, with AUC scores slightly lower than those achieved by the VQC model.
These findings highlight that while QSVM performs comparably to classical SVM, the VQC shows improved performance in terms of AUC and MCC under specific configurations. Table 13 summarizes the key performance metrics.
@@ 7.1 Conclusion
In the Quantum Support Vector Machine (QSVM), the model leverages the inherent characteristic of support vectors, which allows it to define decision boundaries using only a small number of data points. This feature was reflected in our experimental results, where the best performance was achieved with just 20 training samples—yielding an accuracy of 70%, an Area Under the ROC Curve (AUC) of 0.7071, and a Matthews Correlation Coefficient (MCC) of 0.4141.
On the other hand, the Variational Quantum Classifier (VQC) requires iterative optimization to locate a global minimum, continuously updating its parameters to find the optimal solution. Under our maximum setting of 150 epochs, the VQC achieved its best performance, with an accuracy of 73%, an AUC of 0.7032, and an MCC of 0.1475.
@@ Acknowledgements
This work was financially supported by the National Science and Technology Council (NSTC), R.O.C., under Project No. 113-2115-M-A49-015-MY2, titled ”Investigating Quantum Channels in Noisy Quantum Systems: Some Mathematical Problems in Quantum Information Theory.” The authors also acknowledge the assistance of AI-powered language tools in improving the clarity and fluency of the English writing in this manuscript.

### v2

# Comparative Analysis of Quantum Support Vector Machines and Variational Quantum Classifiers for B-cell Epitope Prediction in Vaccine Design

@@ Abstract
Quantum computing offers new opportunities for addressing complex classification tasks in biomedical applications. This study investigates two quantum machine learning models-the Quantum Support Vector Machine (QSVM) and the Variational Quantum Classifier (VQC)-in the context of B-cell epitope prediction, a key step in modern vaccine design. QSVM builds upon the classical SVM framework by using quantum circuits to encode nonlinear kernel computations, while VQC replaces the entire classification pipeline with trainable quantum circuits optimized variationally.
A benchmark dataset from the Immune Epitope Database (IEDB) is used for model evaluation. Each epitope is represented by 10 physicochemical features, and dimensionality reduction via Principal Component Analysis (PCA) is applied to assess model performance across different feature spaces. We also examine the effect of sample size on prediction outcomes.
Experimental results show that QSVM performs well under limited data conditions, while VQC achieves higher accuracy in larger datasets. These findings highlight the potential of quantum-enhanced models for bioinformatics tasks, particularly in supporting efficient and scalable epitope-based vaccine development.
@@ 1 Introduction
The identification of B-cell epitopes—regions of an antigen recognized by the immune system—is a central task in vaccine design, diagnostic development, and therapeutic antibody production. Since the first predictive method was proposed in 1981 [1], linear B-cell epitope prediction has remained a major research focus in computational immunology due to its potential to reduce the cost and time of vaccine development, especially during pandemics such as COVID-19.
A variety of machine learning models have been developed over the past two decades for B-cell epitope prediction. Early tools such as BcePred [2] employed physicochemical propensity scales (e.g., hydrophilicity, flexibility), while BepiPred [3] introduced hidden Markov models (HMMs). Subsequent tools like ABCpred [4] applied recurrent neural networks (RNNs), and methods such as COBEpro [5], SVMTriP [6], and LBEEP [7] leveraged support vector machines (SVMs) alongside novel feature engineering strategies.
Despite these developments, recent benchmarking studies reveal that many of these models exhibit limited performance on independent test datasets. For example, Melagraki et al. [8] compared widely-used predictors and found that BepiPred had the highest Matthews correlation coefficient (MCC) of 0.0778 among per-peptide methods, while other methods such as BcePred, SVMTriP, and LBEEP performed significantly worse (MCC ¡ 0.03). Even consensus ensemble methods showed only modest improvements in overall prediction accuracy. These findings suggest that further methodological innovation is necessary to improve generalization and predictive power for epitope classification tasks.
Quantum machine learning (QML) has emerged as a promising paradigm that combines the power of quantum computation with the flexibility of machine learning frameworks. Two notable QML models are the quantum support vector machine (QSVM) [9] and the variational quantum classifier (VQC) [10], both of which exploit high-dimensional Hilbert spaces and entanglement for learning complex patterns in data. Recent biomedical applications have demonstrated the potential of QML in protein structure analysis, peptide classification, and drug discovery [11, 12].
In this study, we apply QSVM and VQC to the task of linear B-cell epitope prediction. Using benchmark datasets curated from the Immune Epitope Database (IEDB), we preprocess features via principal component analysis (PCA) and evaluate classification performance under varying feature dimensionalities and sample sizes. We further analyze quantum circuit complexity and discuss practical implications for deploying these models on near-term quantum hardware. Our contributions include a side-by-side evaluation of quantum classifiers in a biomedically relevant setting and operational insights into model selection for resource-constrained quantum devices.
@@ 2 Support Vector Machine
Support Vector Machines (SVMs) have long been regarded as one of the most effective classical machine learning algorithms for binary classification tasks. In the context of B-cell epitope prediction, SVM-based models have been widely applied due to their strong theoretical foundations and ability to handle high-dimensional feature spaces. This section introduces the core concepts of SVMs, including their geometric intuition, kernel-based extensions for nonlinear data, and their application in bioinformatics domains.
@@ 3 Background and Data Preprocessing
Figure 1: 2D PCA projection of the input features (adapted from [14]).
Figure 2: 3D PCA projection of the input features (adapted from [14]).
@@ 3.1 B-cell Epitope Prediction Background
B-cells play a central role in the adaptive immune system by recognizing antigenic determinants (epitopes) on protein surfaces and producing antibodies specific to those regions. Accurate identification of linear B-cell epitopes is essential for rational vaccine design, as it enables the selection of high-immunogenicity protein segments for epitope-based vaccines.
In order to address the need for efficient and scalable epitope prediction, especially in the context of emergent viral threats such as SARS-CoV-2, immunoinformatics approaches have become increasingly prevalent. Databases such as the Immune Epitope Database (IEDB) and UniProt provide well-annotated repositories of peptide sequences and their associated immune activities. In our study, we focus on linear B-cell epitope data extracted from IEDB, specifically those associated with IgG-class antibodies, which represent the majority of curated entries.
@@ 4 Quantum Support Vector Machine, QSVM
Quantum Support Vector Machines (QSVMs) extend the classical SVM framework into the quantum computing domain by leveraging quantum-enhanced feature spaces. Instead of relying on explicitly defined kernel functions, QSVMs use parameterized quantum circuits to encode data implicitly into a high-dimensional Hilbert space. These quantum feature maps allow the model to exploit quantum parallelism and entanglement to capture complex correlations that may be inaccessible to classical models.
In this section, we present the theoretical motivation and architectural configuration of QSVMs, highlighting their kernel-based classification mechanism and circuit-level implementation. A schematic of the QSVM workflow is shown in Figure 3.
Figure 3: Configuration diagram of the Quantum Support Vector Machine (adapted from [14]).
@@ 4.1 Kernel-based Classification Using Quantum Circuits
$K(x_{i},x_{j})=|\langle\phi(x_{i})|\phi(x_{j})\rangle|^{2}\,.$
The quantum feature maps $\phi(x)$ are implemented by parametrized quantum circuits. The resulting kernel matrix is estimated by repeatedly executing the circuit and measuring it on a computational basis. The number of times the all-zero outcome is observed determines the kernel estimate. This quantum kernel is then passed to a classical SVM solver for final classification, forming a hybrid quantum-classical architecture.
@@ 4.2 Data Encoding and Feature Mapping
$|x\rangle=U_{\phi}(x)|0\rangle^{\otimes n}\,,$
@@ 5 Variational Quantum Classifier, VQC
Variational Quantum Classifiers (VQCs) represent a different paradigm from kernel-based QSVMs by using trainable parameterized quantum circuits as part of an end-to-end learning model. Unlike QSVMs, which rely on fixed quantum feature maps and classical optimization, VQCs integrate both the feature encoding and decision boundary generation into the same quantum circuit architecture, optimized through a variational approach.
In this section, we describe the conceptual framework and architecture of VQCs, which employ alternating layers of quantum gates with tunable parameters to learn from labeled data. Figure 4 illustrates the high-level structure of a typical VQC setup.
Figure 4: Configuration diagram of the Variational Quantum Classifier (adapted from [14]).
@@ 6 Performance Evaluation
This section presents a comparative performance evaluation of the QSVM and VQC models introduced earlier. The objective is to assess their classification accuracy, training behavior, and sensitivity to feature dimensionality and sample size when applied to B-cell epitope prediction tasks. To ensure a fair comparison, both models were tested using identical preprocessed datasets derived from IEDB, with consistent cross-validation settings and model-independent feature compression through principal component analysis (PCA).
The results are analyzed from both predictive and operational perspectives. In addition to standard performance metrics such as accuracy and precision, we examine quantum-specific aspects such as circuit depth, execution time, and trainability to highlight the practical implications of deploying these models on near-term quantum devices.
@@ 6.2 QSVM
Table 2 shows the results of the QSVM. The highest accuracy, $70\%$, was achieved when using 20 samples with 2-dimensional input features.
@@
Table 1: QSVM Accuracy (adapted from [14]).
@@
Table 2: QSVM AUC value (adapted from [14]).
@@
Table 3: QSVM MCC value (adapted from [14]).
@@ 6.3 The Variational Quantum Classifier
Table 5 presents the accuracy of the VQC with 10 training epochs, where the sample size ranges from 1000 to 4000 and the input dimension varies from 2 to 10.
@@
Table 4: VQC 10 epoch Accuracy (adapted from [14]).
As shown in Table 5, the accuracy generally hovers around $73\%$, with the highest accuracy achieved when the sample size is 1000 and the input dimension is 4.
Table 6 presents the AUC values for 10 training epochs, with sample sizes ranging from 1000 to 4000 and input dimensions from 2 to 10.
@@
Table 5: VQC 10 epoch AUC value (adapted from [14]).
As shown in Table 6, the best AUC value of 0.6750 was obtained when the sample size was 4000 and the input dimension was 10.
Table 7 presents the MCC values for 10 training epochs, with sample sizes ranging from 1000 to 4000 and input dimensions from 2 to 10.
@@
Table 6: VQC 10 epoch MCC value (adapted from [14]).
In Table 7, the highest MCC value of 0.0987 was achieved when the sample size was 3000 and the input dimension was 4.
Table 8 presents the accuracy results for 100 training epochs, with sample sizes ranging from 1000 to 4000 and input dimensions from 2 to 10.
@@
Table 7: VQC 100 epoch Accuracy (adapted from [14]).
The accuracy values in Table 8 are similar to those observed at 10 epochs, remaining around $73\%$. In this case, the highest accuracy was achieved with a sample size of 3000 and an input dimension of 10.
Table 9 presents the AUC values for 100 training epochs, with sample sizes ranging from 1000 to 4000 and input dimensions from 2 to 10.
@@
Table 8: VQC 100 epoch AUC value (adapted from [14])
In Table 9, the AUC value reaches 0.7, which occurs at an input dimension of 10 with a sample size of 4000—the largest sample size and the highest dimensional setting in the experiment.
Table 10 presents the MCC values for 100 training epochs, with sample sizes ranging from 1000 to 4000 and input dimensions from 2 to 10.
@@
Table 9: VQC 100 epoch MCC value (adapted from [14]).
As shown in Table 10, the MCC value exceeds 0.1 when the sample size is 1000 and the input dimension is 6, representing a relatively good performance.
Table 11 presents the accuracy results for 150 training epochs, with sample sizes ranging from 1000 to 4000 and input dimensions from 2 to 10.
@@
Table 10: VQC 150 epoch Accuracy (adapted from [14]).
Compared to the previous two epoch settings (Tables 5 and 8), Table 11 shows a slight increase in accuracy, reaching nearly $74\%$.
Table 12 presents the AUC values for 150 training epochs, with sample sizes ranging from 1000 to 4000 and input dimensions from 2 to 10.
@@
Table 11: VQC 150 epoch AUC value (adapted from [14]).
@@
Table 13 presents the MCC values for 150 training epochs, with sample sizes ranging from 1000 to 4000 and input dimensions from 2 to 10.
@@
Table 12: VQC 150 epoch MCC value (adapted from [14]).
@@ 6.4 Effect of Feature Dimensionality
To investigate the influence of feature dimensionality on model performance, we conducted a series of experiments using Principal Component Analysis (PCA) to reduce input features to 2 through 10 components. QSVM and VQC were evaluated under these conditions using fixed training sizes (e.g., 1000, 2000, 4000 samples).
Our results indicate that model performance varies non-linearly with feature dimensionality. In particular, the VQC performs better in low-to-medium dimensional regimes (e.g., 4–6 dimensions), whereas the QSVM is more stable across all tested dimensions. These observations suggest VQC may benefit from feature compression under limited quantum hardware resources.
@@ 6.5 Comparison with Classical Models
We trained a classical Support Vector Machine (SVM) using the radial basis function (RBF) kernel on the same dataset and preprocessing pipeline to establish a baseline. The classical SVM achieved an accuracy of approximately 70%, with AUC scores slightly lower than those achieved by the VQC model.
These findings highlight that while QSVM performs comparably to classical SVM, the VQC performs better in terms of AUC and MCC under specific configurations. Table 13 summarizes the key performance metrics.
@@ 7 Conclusion and Future Work
This final section summarizes the main findings of our comparative study on quantum-enhanced machine learning methods for B-cell epitope prediction. By evaluating QSVM and VQC on benchmark datasets under varying conditions, we aimed to understand the strengths and limitations of each quantum model in both predictive and operational terms. The following conclusions highlight the key insights obtained from this work and offer guidance for future exploration and deployment of quantum classifiers in bioinformatics.
@@ 7.1 Conclusion
In the Quantum Support Vector Machine (QSVM), the model leverages the inherent characteristic of support vectors, which allows it to define decision boundaries using only a small number of data points. This feature was reflected in our experimental results, where the best performance was achieved with just 20 training samples—yielding an accuracy of $70\%$, an Area Under the ROC Curve (AUC) of 0.7071, and a Matthews Correlation Coefficient (MCC) of 0.4141.
On the other hand, the Variational Quantum Classifier (VQC) requires iterative optimization to locate a global minimum, continuously updating its parameters to find the optimal solution. Under our maximum setting of 150 epochs, the VQC achieved its best performance, with an accuracy of $73\%$, an AUC of 0.7032, and an MCC of 0.1475.
@@ Acknowledgements
This work was financially supported by the National Science and Technology Council (NSTC), Taiwan, under Project No. 113-2115-M-A49-015-MY2, titled ”Investigating Quantum Channels in Noisy Quantum Systems: Some Mathematical Problems in Quantum Information Theory.” The authors also acknowledge the assistance of AI-powered language tools in improving the clarity and fluency of the English writing in this manuscript.